# A Queueing Theoretic Perspective on Low-Latency LLM Inference with Variable Token Length

### v1

@@ I Introduction
A Large Language Model (LLM) is a gigantic neural network trained on massive amount of text data such as GPT [1] and LLaMA [2]. It is not only capable of generating natural language sentences, but also possesses the power of understanding textual meaning. Nowadays, LLMs have been comprehensively applied in almost every aspect of content generation, and rapidly expand to search engine and software engineering. It is even believed that LLMs even light the way toward artificial general intelligence.
As a generative model, the way that a LLM creates content is called “inference”. The LLM resides at one or more computing nodes, and users submit their inference requests to the LLM platform for processing. On the surface, LLM inference is nothing different from traditional cloud computing service, while the Transformer structure inside LLM models has made this process unique. As the core of Transformer, the attention mechanism is able to capture the long-range dependency across data by executing a softmax function along with matrix products [3]. This operation consists of the “prefill” and “decode” stages in LLM inference. The input tokens of all requests are computed and the corresponding intermediate results are cached at the key-value (KV) store [4] in the prefill stage. The output token is generated one-by-one autoregressively in the decode stage. In terms of resource usage, prefill is compute intensive, while decode is memory access intensive [5]. Intuitively, an input or output request with larger token length demands more time to prefill and decode, thus affecting the latency of LLM inference.
@@ II-B Impact of Tokens on Inference latency
Both input and output tokens must undergo the inference and computation processes of large language models (LLMs), impacting the model’s inference latency. We represent the workload of an inference task as the number of tokens to be generated, and measure the inference latency with regard to different numbers of tokens. Consider that a LLaMA-2-7b-chat model runs on an NVIDIA A100 GPU. Table I shows the inference latency when the pair of input and output tokens change.
@@ IV-B LLMs Dynamic Batching
Distinguished from traditional non-AIGC models, the inference time of a LLM request is not deterministic, but depending on the numbers of input and output tokens. Within a batch, different requests may have varying numbers of input (and output) tokens, necessitating padding to align with the longest input (and output). Consequently, the inference time of a batch is influenced not only by the batch size but also by the maximum input and output token sizes. For requests involving long text inputs, such as article translation and summarization, users are typically less sensitive to queuing delays. Therefore, our focus here is primarily on short text inputs. We suppose the number of input tokens follows the uniform distribution from 00 to $m_{1}$ ($m_{1}\geq 0$). For a batch $b$, the CDF of the max input token size $D_{max}$ is given by:
@@ IV-B1 Uniform Distribution of Output Tokens
we suppose that the number of output tokens $N$ follows the uniform distribution from 00 to $m_{2}$ ($m_{2}\geq 0$).
Let $L$ be a generic random variable of the max output token size in the batch, so that $E[L]=\frac{m_{2}b}{b+1}$ For a batch with size $b$, the mean inference time is given by:
@@ IV-B2 Truncated Gaussian Distribution of Output Tokens
[…]
@@ IV-D LLMs Elastic Batching
$\begin{split}H^{[b,n_{b}]}&{=}k_{1}b{+}k_{2}{+}(k_{3}b{+}k_{4})n_{1}{+}(k_{3}(b{-}1){+}k_{4})(n_{2}{-}n_{1})\\ &{+}\cdots{+}(k_{3}1(b{-}(b{-}1)){+}k_{4})(n_{b}{-}n_{b-1})\\ &{=}k_{1}b{+}k_{2}{+}k_{3}(n_{1}{+}n_{2}{+}\cdots{+}n_{b}){+}k_{4}n_{b}\\ &{=}k_{1}b{+}k_{2}{+}k_{3}bE[N]{+}k_{4}n_{b}.\end{split}$ (26)
@@ V-A Experiment Setup
In this section, we analyze the inference latency curve based on output token size and batch size using a real LLM-based inference server on NVIDIA A100. Subsequently, we validate the practicality of our derived mathematical expressions through numerical experiments conducted using MATLAB’s event-driven simulation. Our experiments include three instruction-finetuned language models: LLaMA-2-7b-chat, Chinese-Alpaca-2-7b, and ChatGLM3-6b. For experimental data, we simulate virtual scenarios with output token sizes following various distributions.
@@ V-B Impact of Token Limit
We assume that the output token size follows a logarithmic normal distribution with log mean 7 and log standard deviation 0.7. Fig.4a, Fig.4b and Fig.4c show the mean queuing delay $E[W]$, $E[W_{qs}]$ and loss fraction $\pi(\tau)$ with regard to the max token limit $n_{max}$. We randomly set the arrival rate such that $\rho<1$. The arrival rate $\lambda$ is 1/40 in Fig. 4a and $\lambda$ is 1/25 and $\tau$ is 60 in Fig. 4b and Fig.4c. We observed that each curve is accurately described by the provided expressions. When assuming all users are sufficiently patient, the mean queuing delay tends to increase with the maximum token limit. This is because a significant number of users requiring long responses enter the system, significantly slowing down LLMs inference. However, when considering user impatience, the mean queuing delay gradually stabilizes. This occurs because more users leave the system without being served, reducing the overall load over time.
@@
(a) $\lambda=0.43$
(b) Dynamic batching with max batch size and elastic batching

### v2

@@ I Introduction
A Large Language Model (LLM) is a gigantic neural network trained on massive amount of text data such as GPT [37] and LLaMA [3]. It is not only capable of generating natural language sentences, but also possesses the power of understanding textual meaning. Nowadays, LLMs have been comprehensively applied in almost every aspect of content generation, and rapidly expand to search engine and software engineering. It is even believed that LLMs even light the way toward artificial general intelligence. As a generative model, the way that a LLM creates content is called “inference”. The LLM resides at one or more computing nodes, and users submit their inference requests to the LLM platform for processing. Intuitively, an input or output request with larger token length demands more time, thus affecting the latency of LLM inference.
@@ II-B Impact of Tokens on Inference latency
Both input and output tokens must undergo the inference and computation processes of large language models (LLMs), impacting the model’s inference latency. We represent the workload of an inference task as the number of tokens to be generated, and measure the inference latency with regard to different numbers of tokens. Consider that a LLaMA-2-7b-chat model runs on an NVIDIA A100 GPU. Table I shows the inference latency when the pair of input and output tokens change. In each experiment, we collect 100 conversation requests with the LLM, record the generated inference latency, and compute their average values.
@@ IV-B LLMs Dynamic Batching
Distinguished from traditional non-AIGC models, the inference time of a LLM request is not deterministic, but depending on the numbers of input and output tokens. Within a batch, different requests may have varying numbers of input (and output) tokens, necessitating padding to align with the longest input (and output). Consequently, the inference time of a batch is influenced not only by the batch size but also by the maximum input and output token sizes. For requests involving long text inputs, such as article translation and summarization, users are typically less sensitive to queuing delays. Therefore, our focus here is primarily on short text inputs. We suppose the number of input tokens follows the uniform distribution from 0 to $m_{1}$ ($m_{1}\geq 0$). For a batch $b$, the CDF of the max input token size $D_{max}$ is given by:
@@ IV-B1 Uniform Distribution of Output Tokens
we suppose that the number of output tokens $N$ follows the uniform distribution from 0 to $m_{2}$ ($m_{2}\geq 0$).
Let $L$ be a generic random variable of the max output token size in the batch, so that $E[L]=\frac{m_{2}b}{b+1}$ . For a batch with size $b$, the mean inference time is given by:
@@ IV-B2 Truncated Gaussian Distribution of Output Tokens
$F_{N_{tr}}(x)$ $=$ $\int_{0}^{x}\frac{\phi(\frac{t-u}{\sigma})}{\sigma(1-\Phi(\frac{-u}{\sigma}))}dt,x\geq 0,$ (22)
@@ IV-D LLMs Elastic Batching
$\begin{split}H^{[b,n_{b}]}&{=}k_{1}b{+}k_{2}{+}(k_{3}b{+}k_{4})n_{1}{+}(k_{3}(b{-}1){+}k_{4})(n_{2}{-}n_{1})\\ &{+}\cdots{+}(k_{3}1(b{-}(b{-}1)){+}k_{4})(n_{b}{-}n_{b-1})\\ &{=}k_{1}b{+}k_{2}{+}k_{3}bE[N]{+}k_{4}n_{b}.\end{split}$ (26)
@@ V-A Experiment Setup
In this section, we analyze the inference latency curve based on output token size and batch size using a real LLM-based inference server on NVIDIA A100. Subsequently, we validate the practicality of our derived mathematical expressions through numerical experiments conducted using MATLAB’s event-driven simulation. The base language models can only perform text continuation, i.e., generating the following text when given the preceding context, which cannot engage in conversational interaction. In contrast, an instruction-finetuned language models can understand user requests and provide appropriate responses. Therefore, three types of open-source instruction-finetuned language models are used in this experiment: LLaMA-2-7b-chat, Chinese-Alpaca-2-7b, and ChatGLM3-6b. For experimental data, we simulate virtual scenarios with output token sizes following various distributions.
@@ V-B Impact of Token Limit
We assume that the output token size follows a logarithmic normal distribution with log mean 7 and log standard deviation 0.7. Fig.4a, Fig.4b and Fig.4c show the mean queuing delay $E[W]$, $E[W_{qs}]$ and loss fraction $\pi(\tau)$ with regard to the max token limit $n_{max}$. Among them, the red, blue, and black lines represent the numerical results obtained through mathematical equations, while the red triangles, blue squares, and black circles correspond to the simulation results. We randomly set the arrival rate such that $\rho<1$. The arrival rate $\lambda$ is 1/40 in Fig. 4a and $\lambda$ is 1/25 and $\tau$ is 60 in Fig. 4b and Fig.4c. We observed that each curve is accurately described by the provided expressions. When assuming all users are sufficiently patient, the mean queuing delay tends to increase with the maximum token limit. This is because a significant number of users requiring long responses enter the system, significantly slowing down LLMs inference. However, when considering user impatience, the mean queuing delay gradually stabilizes. This occurs because more users leave the system without being served, reducing the overall load over time.
@@ V-C Impact of Batching Strategies
(a) $\lambda=0.43$
(b) Dynamic batching with max batch size and elastic batching